Grade 12Algebra

. If the product of the roots of the equation, x 2−5kx+2e2loge ?k?−1=0 is 49, then the sum of the squares of the roots of the equation is :

azad
7 Years agoGrade 12

1 Answer

Aditya Gupta
7 Years ago
The question is not clear because of question marks in between the equations. Could you repost the question along with a picture so I can assist you further? Btw the sum of squares of roots can be written as (sum of roots)^2 - 2*product of roots. We know that sum of roots= -b/a and product= c/a